OpenStudy (anonymous):

53 krissywatts find the quadratic function y=ax^2+bx+c whose graph passes through the given points. (-1,14) (1,4) (3,2)

Directrix (directrix):

Recall that these point represent x and y-coordinate relationships on the curve. Since the curve passes through (-1,14) , we obtain that: y = ax^2 + bx + c ==> 14 = a(-1)^2 + b(-1) + c ==> a - b + c = 14 Since it passes through (1,4) we obtain: y = ax^2 + bx + c ==> 4 = a(1)^2 + b(1) + c ==> a + b + c = 4 Then, it passes through (3,2) to get: 2 = a(3)^2 + b(3) + c ==> 9a + 3b + c = 2 Thus creating this system of equations: { a - b + c = 14 { a + b + c = 4 { 9a + 3b + c = 2 This system has solutions a = 1, b = -5, and c = 8. Therefore, the quadratic function is y = x^2 - 5x + 8. -------- Use the online calculator to solve 3 equations in 3 unknowns at the link: http://www.analyzemath.com/Calculators/Calculator_syst_eq.html

OpenStudy (anonymous):

Another approach to the solution is attached.